# Survey of Data-driven Newsvendor: Unified Analysis and Spectrum of Achievable Regrets

### v4

@@ Abstract
In the Newsvendor problem, the goal is to guess the number that will be drawn from some distribution, with asymmetric consequences for guessing too high vs. too low. In the data-driven version, the distribution is unknown, and one must work with samples from the distribution. Data-driven Newsvendor has been studied under many variants: additive vs. multiplicative regret, high probability vs. expectation bounds, and different distribution classes. This paper studies all combinations of these variants, filling many gaps in the literature and simplifying many proofs. In particular, we provide a unified analysis based on a notion of clustered distributions, which in conjunction with our new lower bounds, shows that the entire spectrum of regrets between $1/\sqrt{n}$ and $1/n$ is possible. Simulations on commonly-used distributions demonstrate that our notion is the ?correct? predictor of empirical regret across varying data sizes.
@@ 1.1 Existing and New Results
Upper Bound $\beta=0,\infty$ known (lin2022data) $\beta=\infty$ known (besbes2023big) 333besbes2023big attribute this result to levi2015data, but to the best of our understanding, levi2015data is insufficient because its proof only holds for $\epsilon\leq 1$. Therefore, we attribute this result to besbes2023big instead.
@@ 1.2 Further Related Work
Sample complexity has roots in statistical learning theory, which typically studies classification and regression problems under restricted hypothesis classes (shalev-shwartz2014understanding; mohri2018foundations). Its concepts can also be extended to general decision problems (balcan2020data; balcan2021much), or even specific inventory policy classes (xie2024vc). However, data-driven Newsvendor results differ by considering multiplicative regret, having a specialized but unbounded loss function (there are no assumptions on demand being bounded), and typically requiring analyses that are tighter than uniform convergence. In data-driven Newsvendor, it is also difficult to directly convert high-probability bounds into expectation bounds via an integral (see e.g. besbes2023big, Lem. E-5), because the regret can be unbounded, while high-probability bounds only hold for small values of $\varepsilon$ (or equivalently, large values of $n$). Our results further differ by considering specific restrictions on the distribution $F$.
@@
Big-data Newsvendor is a generalization of data-driven Newsvendor where past demand samples are accompanied by contextual information, and the inventory decision can be made knowing the future context. This model was popularized by ban2019big, and motivated by the notion of contexts from machine learning. Meanwhile, data-driven inventory is a generalization of data-driven Newsvendor where one is re-stocking a durable good over multiple periods, that was also considered in the original paper by levi2007provably. Further variants of this model include learning censored demands (huh2009nonparametric), capacitated order sizes (cheung2019sampling), lost sales (zhang2020closing), and pricing (chen2021nonparametric). Our paper focuses on a single period without contexts, and does not aim to cover these generalizations.
@@
Meanwhile, other works impose alternative assumptions on the underlying distribution to achieve similar faster rates for SAA on data-driven Newsvendor. An example is the Increasing Failure Rate (IFR) property, which requires 1F1-F to be log-concave. Under this assumption, and the assumption that $F$ is a continuous distribution, zhang_sampling-based_2025 establishes a sample complexity of $O\left((1+\varepsilon^{-1/2}+\varepsilon^{-1})\log(1/\delta)\right)$, where $\varepsilon$ is the multiplicative regret. When $\varepsilon$ is close to 0, this result implies that the high probability multiplicative regret is $O\left(\frac{\log(1/\delta)}{n}\right)$, which is the same as our result for clustered distributions with $\beta=0$. In fact, we show in Appendix C that any continuous distribution with the IFR property is $(0,\gamma,\zeta)$-clustered for some $\gamma$ and $\zeta$, and therefore our result for $\beta=0$ can be viewed as a generalization of their result.
@@ 6 Simulations
We numerically compute the expected additive regrets for several distributions under different values of $q$ and number of samples $n$. Additionally, we provide the 95th percentile of additive regret in Appendix E, which represents the high probability additive regret with $\delta=0.05$. By comparing these simulation results with the relative order of $\beta$ across distributions, we show that our theory largely captures the comparison of regrets across distributions as the number of samples grows. In particular, it helps explain the ?crossover points? in the regret curves, a phenomenon that previous theories relying on lower-bounding the PDF (besbes2013implications; lin2022data) fail to capture.
@@ Appendix A Examples of Clustered Distributions
Bernoulli $c\cdot\mathrm{Ber}(p)$
$F(a)=\begin{cases}0,&a<0\\ 1-p,&a\in[0,c)\\ 1,&a\geq c\end{cases}$
$\begin{cases}0,&\text{if }\frac{1}{\sqrt{n}}<|1-p-q|\\ &\text{or }\frac{1}{\sqrt{n}}=1-p-q\\ c,&\text{otherwise}\end{cases}$
$\begin{cases}\infty,&\text{if }p+q=1\\ &\text{or }\gamma\zeta=1\\ \frac{\ln|1-p-q|}{\ln\gamma+\ln c}-1,&\text{if }p+q\neq 1\\ &\text{and }\gamma\zeta\in(\frac{1}{\sqrt{n}},1)\\ 0,&\text{otherwise}\end{cases}$
@@
$F(a)=\frac{a-b_{1}}{b_{2}-b_{1}}$
$\forall a\in[b_{1},b_{2}]$
$\begin{cases}\frac{b_{2}-b_{1}}{\sqrt{n}},&\text{if }\frac{1}{\sqrt{n}}\leq\max\{q,1-q\}\\ \makebox[0.0pt][l]{$(b_{2}-b_{1})\max\{q,1-q\}$,}\\ &\text{otherwise}\end{cases}$
$\begin{cases}\infty,&\text{if }\gamma\zeta=1\\ -\frac{\ln n}{2(\ln\gamma+\ln\zeta)}-1,&\text{if }\gamma\zeta\in(\frac{1}{\sqrt{n}},1)\\ 0,&\text{otherwise}\end{cases}$
@@
$F(a)=1-e^{-\lambda a}$
$\forall a>0$
$\begin{cases}\frac{1}{\lambda}\ln\frac{1-q}{1-q-\frac{1}{\sqrt{n}}},&\text{if }\frac{1}{\sqrt{n}}<1-q\\ \infty,&\text{otherwise}\end{cases}$
@@
$F(a)=1-\left(\frac{x_{m}}{a}\right)^{\alpha}$
$\forall a\geq x_{m}$
$\begin{cases}\makebox[0.0pt][l]{$\frac{x_{m}}{(1-q-\frac{1}{\sqrt{n}})^{\frac{1}{\alpha}}}-\frac{x_{m}}{(1-q)^{\frac{1}{\alpha}}}$,}\\ &\text{if }\frac{1}{\sqrt{n}}<1-q\\ \infty,&\text{otherwise}\end{cases}$
@@
Based on Table 2, we compute the exact values of $\zeta$ and $\beta$ for all distributions considered in Section 6. The results are reported in Table 3. Consistent with the setup in Section 6, we report values for $q=0.4$ and $q=0.9$. For each value of $q$, we present results for both a small number of samples ($n=11$) and a large number of samples ($n=196$). The values of $n$ are chosen to remain within the range 1 to 200 used in Section 6 and correspond to points where we computed regrets (every 5 integers, starting with 1) and correspond to points where we computed regrets (every 5 integers, starting with 1). The fixed value of $\gamma$ is smaller when $q=0.9$, which reflects the smaller slope at $a^{*}$ for the Exponential, Pareto, and Log-normal distributions.

### v5

@@ Abstract
In the Newsvendor problem, the goal is to guess the number that will be drawn from some distribution, with asymmetric consequences for guessing too high vs. too low. In the data-driven version, the distribution is unknown, and one must work with samples from the distribution. Data-driven Newsvendor has been studied under many variants: additive vs. multiplicative regret, high-probability vs. expectation bounds, and different distribution classes. This paper studies all combinations of these variants, filling many gaps in the literature and simplifying many proofs. In particular, we provide a unified analysis based on a notion of clustered distributions, which in conjunction with our new lower bounds, shows that the entire spectrum of regrets between $1/\sqrt{n}$ and $1/n$ is possible. Simulations on commonly-used distributions demonstrate that our notion is the ?correct? predictor of empirical regret across varying data sizes.
Keywords: data-driven decision-making, Newsvendor, distribution classes, learning theory
@@ 1.1 Existing and New Results
Upper Bound $\beta=0,\infty$ known (lin2022data) $\beta=\infty$ known (besbes2023big) 333besbes2023big attribute this result to levi2015data, but to the best of our understanding, levi2015data is insufficient because its proof only holds for $\epsilon\leq 1$. Therefore, we attribute this result to besbes2023big instead. We note that besbes2023big works if an upper bound on the mean is known and one uses projected SAA instead—see Appendix D.
@@ 1.2 Further Related Work
Sample complexity has roots in statistical learning theory, which typically studies classification and regression problems under restricted hypothesis classes (shalev-shwartz2014understanding; mohri2018foundations). Its concepts can also be extended to general decision problems (balcan2020data; balcan2021much), or even specific inventory policy classes (xie2024vc). However, data-driven Newsvendor results differ for various reasons: considering multiplicative regret instead of only additive regret, having a specialized but unbounded loss function (there are no assumptions on demand being bounded), and typically requiring analyses that are tighter than uniform convergence. In data-driven Newsvendor, it is also difficult to directly convert high-probability bounds into expectation bounds unless one knows an upper bound on the mean, because the regret can be unbounded, while high-probability bounds only hold for small values of $\varepsilon$ or equivalently large values of $n$ (see Appendix D). Our results further differ by considering specific restrictions on the distribution $F$.
@@
Big-data Newsvendor is a generalization of data-driven Newsvendor where past demand samples are accompanied by contextual information, and the decision can be made knowing the future context. This model was popularized by ban2019big, and motivated by the notion of contexts from machine learning. Meanwhile, data-driven inventory is a generalization of data-driven Newsvendor where one is re-stocking a durable good over multiple periods, that was also considered in the original paper by levi2007provably. Further variants include censored demands when sales are lost (e.g. huh2009nonparametric; besbes2013implications; zhang2020closing; hssaine2024data), capacitated order sizes (e.g. cheung2019sampling), and pricing (e.g. chen2021nonparametric; chen2022dynamic; chen2024optimal). Our paper focuses on a single period without contexts, and does not aim to cover these generalizations.
@@
Meanwhile, other works impose alternative assumptions on the underlying distribution to achieve similar faster rates for SAA on data-driven Newsvendor. An example is the Increasing Failure Rate (IFR) property, which requires 1F1-F to be log-concave. Under this assumption, and the assumption that $F$ is a continuous distribution, zhang_sampling-based_2025 establishes a sample complexity of $O\left((1+\varepsilon^{-1/2}+\varepsilon^{-1})\log(1/\delta)\right)$, where $\varepsilon$ is the multiplicative regret. When $\varepsilon$ is close to 0, this result implies that the high-probability multiplicative regret is $O\left(\frac{\log(1/\delta)}{n}\right)$, which is the same as our result for clustered distributions with $\beta=0$. In fact, we show in Appendix C that any continuous distribution with the IFR property is $(0,\gamma,\zeta)$-clustered for some $\gamma$ and $\zeta$, and therefore our result for $\beta=0$ can be viewed as a generalization of their result.
@@ 6 Simulations
We numerically compute the expected additive regrets for several distributions under different values of $q$ and number of samples $n$. Additionally, we provide the 95th percentile of additive regret in Appendix E, which represents the high-probability additive regret with $\delta=0.05$. By comparing these simulation results with the relative order of $\beta$ across distributions, we show that our theory largely captures the comparison of regrets across distributions as the number of samples grows. In particular, it helps explain the ?crossover points? in the regret curves, a phenomenon that previous theories relying on lower-bounding the PDF (besbes2013implications; lin2022data) fail to capture.
@@ Results.
7 Conclusion
We provide a survey of results and (simplified) proofs for data-driven Newsvendor, including both upper and lower bound analyses, varying along several dimensions: additive vs. multiplicative regret, high-probability vs. expectation bounds, and different distribution classes. We introduce a notion of clustered distributions based on the CDF, which shows the entire spectrum of convergence rates between $1/\sqrt{n}$ and $1/n$ is possible, and is also a useful predictor of empirical regret in simulations. We hope this can be a useful reference for future scholars of data-driven Newsvendor, which can be broadly viewed as a foundation for data-driven decision making in Operations Research.
Acknowledgments.
The authors thank Meichun Lin for identifying an earlier error in a comparison to lin2022data, Omar Besbes and Omar Mouchtaki for useful discussions related to the paper, and the anonymous referees and associate editor of Operations Research as well as area editor Dennis Zhang for their comprehensive comments that improved the paper.
@@ Appendix A Examples of Clustered Distributions
Bernoulli $c\cdot\mathrm{Ber}(p)$ $F(a)=\begin{cases}0,&a<0\\ 1-p,&a\in[0,c)\\ 1,&a\geq c\end{cases}$ $\begin{cases}0,&\text{if }\frac{1}{\sqrt{n}}<|1-p-q|\\ &\text{or }\frac{1}{\sqrt{n}}=1-p-q\\ c,&\text{otherwise}\end{cases}$ $\begin{cases}\infty,&\text{if }p+q=1\\ &\text{or }\gamma\zeta=1\\ \frac{\ln|1-p-q|}{\ln\gamma+\ln c}-1,&\text{if }p+q\neq 1\\ &\text{and }\gamma\zeta\in(\frac{1}{\sqrt{n}},1)\\ 0,&\text{otherwise}\end{cases}$
@@
Based on Table 2, we compute the exact values of $\zeta$ and $\beta$ for all distributions considered in Section 6. The results are reported in Table 3. Consistent with the setup in Section 6, we report values for $q=0.4$ and $q=0.9$. For each value of $q$, we present results for both a small number of samples ($n=11$) and a large number of samples ($n=196$). The values of $n$ are chosen to remain within the range 1 to 200 used in Section 6 and correspond to points where we computed regrets (every 5 integers, starting with 1). The fixed value of $\gamma$ is smaller when $q=0.9$, which reflects the smaller slope at $a^{*}$ for the Exponential, Pareto, and Log-normal distributions.